# NLPrompt: Noise-Label Prompt Learning for Vision-Language Models

### v1

@@ Abstract
The emergence of vision-language foundation models, such as CLIP, has revolutionized image-text representation, enabling a broad range of applications via prompt learning. Despite its promise, real-world datasets often contain noisy labels that can degrade prompt learning performance. In this paper, we demonstrate that using mean absolute error (MAE) loss in prompt learning, named PromptMAE, significantly enhances robustness against noisy labels while maintaining high accuracy. Though MAE is straightforward and recognized for its robustness, it is rarely used in noisy-label learning due to its slow convergence and poor performance outside prompt learning scenarios. To elucidate the robustness of PromptMAE, we leverage feature learning theory to show that MAE can suppress the influence of noisy samples, thereby improving the signal-to-noise ratio and enhancing overall robustness. Additionally, we introduce PromptOT, a prompt-based optimal transport data purification method to enhance the robustness further. PromptOT employs text encoder representations in vision-language models as prototypes to construct an optimal transportation matrix. This matrix effectively partitions datasets into clean and noisy subsets, allowing for the application of cross-entropy loss to the clean subset and MAE loss to the noisy subset. Our Noise-Label Prompt Learning method, named NLPrompt, offers a simple and efficient approach that leverages the expressive representation and precise alignment capabilities of vision-language models for robust prompt learning. We validate NLPrompt through extensive experiments across various noise settings, demonstrating significant performance improvements.
@@ 1 Introduction
In the realm of noisy label learning, mean absolute error (MAE) has been identified as a robust loss function within the traditional training paradigm [17]. However, MAE often suffers from slow convergence and poor performance during training, making it seldom employed as a classification loss in noise-label learning. Nevertheless, our investigation reveals an interesting phenomenon: employing MAE loss in Prompt learning (PrompMAE) notably enhances robustness while maintaining high accuracy compared to traditional cross-entropy loss. As demonstrated in Figure 1, MAE exhibits strong accuracy and fast convergence even in the presence of substantial noise.
@@
To elucidate the robustness of PromptMAE, we leverage feature learning theory [1, 3, 36], which categorizes latent representations into task-relevant and task-irrelevant components. By analyzing the optimization dynamics of these features with gradient-descent-based training, we can gain valuable insights into convergence and generalization. To this end, we find that robust prompt learning is achieved when task-relevant features dominate. Our analysis indicates that PrompMAE can suppress the influence of noisy samples, thereby enhancing robustness in prompt learning for vision-language models.
@@
In this paper, we introduce PromptOT, a prompt-based optimal transport data purification method, designed to enhance the robustness of prompt learning in vision-language foundation models. PromptOT leverages the text encoder as prototypes for the transportation matrix, facilitating robust prompt learning by partitioning the dataset into clean and noisy subsets. Recognizing that cross-entropy (CE) loss generally outperforms MAE on clean datasets, we apply MAE loss to train the noisy subset and cross-entropy loss to train the clean subset. This dual strategy, supported by PromptOT purification, harmonizes the strengths of both MAE and CE loss under varying noisy conditions. Our comprehensive method, named NLPrompt, leverages the expressive representation and alignment capabilities of vision-language models, offering a simple and efficient solution for robust prompt learning in the presence of noisy labels. In summary, our contributions are threefold:
@@ 2.3 Feature Learning Theory
To further understand how noisy label learning affects prompt learning, we leverage feature learning theory to analyze the learning process. Feature learning theory [1, 3, 23] categorizes latent features into task-relevant and task-irrelevant components, expressing the trainable weights as a combination of these feature types. From this perspective, feature learning analyzes the coefficients of these features to gain insight into learning dynamics. Beyond its application in traditional learning paradigms, prompt learning can also be explained by feature learning theory [36]. In this paper, we adopt feature learning theory to demonstrate the robustness of MAE in prompt learning.
@@ 3 Preliminary
$\min\limits_{\mathbf{Q}\in\mathbb{R}_{+}^{n\times m}}\qquad$ $\langle\mathbf{C},\mathbf{Q}\rangle$ (1)
s.t. $\mathbf{Q}\mathds{1}_{m}=\bm{\alpha},\ \mathbf{Q}^{\top}\mathds{1}_{n}=\bm{\beta}$ (2)
@@
s.t. $\mathbf{Q}\mathds{1}_{m}=\bm{\alpha},\ \mathbf{Q}^{\top}\mathds{1}_{n}=\bm{\beta}$
@@ 5 Methodology
Table 1: Performance metrics across various datasets and noise levels.
@@ 6.3 Implementation Details
In our experiments, we adopt the same setup as CoOp [55] and JoAPR [19] to ensure a fair comparison. We use the SGD optimizer with an initial learning rate of 0.002 and employ cosine annealing. Our model backbone is consistent with CoOp [55], based on the pre-trained CLIP model [42], utilizing either ResNet-50 or ViT-B/16 as the image encoder, with ResNet-50 as the default if not explicitly specified. We use a 63M parameter text transformer as the text encoder. The default number of training epochs is set to 200. Additionally, we employ 16 shared context tokens across all categories, with the class token placed at the end. We sample a 16-shot training set from each dataset and evaluate the model on the original test set. The reported experimental results are the averages of test accuracy from three runs with different seeds, with the highest accuracy highlighted in bold. All experiments were conducted using PyTorch [38] on a cluster equipped with NVIDIA A40 GPU.
@@ 6.4 Performance Comparison
Table 2: Test accuracy (%) on a real-world noisy dataset Food101N.
Method CoOp GCE JoAPR NLPrompt
Accuracy 69.50 71.32 72.57 76.46
6.5 Experiments on Purfication Strategy
In this section, we present experiments to demonstrate that PromptOT is an effective purification method in prompt learning. We frame the purification task as a binary classification problem, where the goal is to distinguish between clean and noisy samples. To evaluate the purification performance, we use accuracy and F1-score metrics for this binary classification task. We compare our purification strategy with the partition method that uses pseudo-labels generated by CLIP zero-shot (denoted as CLIP-ZS), as well as the partition strategy used in JoAPR [19]. For our experiments, we use the Caltech101, DTD, and Flowers datasets, with results shown in Table 3. From this table, we observe that our prompt-based OT selection method achieves higher purification accuracy compared to both the CLIP zero-shot partition and the JoAPR partition.
Table 3: Comparison between different purification strategies. (%)
Method Caltech101 DTD Flowers
Acc F1 Acc F1 Acc F1
CLIP-ZS 70.38 58.35 56.25 26.72 56.62 25.16
JoAPR 50.88 44.99 51.54 32.32 50.37 32.56
NLPrompt 75.37 67.41 59.97 36.89 62.07 39.49
@@ 6.6 Few-shot Learning Analysis
6.7 Ablation Study
To evaluate the effectiveness of each component of our method, we conduct ablation studies on the Flowers102 dataset. Here, we employ ViT-B/16 as the backbone and train for 100 epochs in the symmetric noise scenario. The experimental results are shown in table 4. To validate the effectiveness of OT, we designed two sets of experiments: one without using OT for data purification and another using OT for data purification. The experimental design is as follows: (a) Use CE loss for all data; (b) Use MAE loss for all data; (c) Use random initialization prototype instead of CLIP text feature as initialization; (d) Use CE loss for clean data only after removing noisy data; (e) Use MAE loss for noisy data only after removing clean data. The average results show that (d) outperforms (a), and (e) outperforms (b), further validating the effectiveness of PromptOT in the data purification process. In comparison, our NLPrompt method integrates OT for purification, using CE loss for clean data and MAE loss for noisy data. This approach not only benefits from OT but also optimizes different data types with appropriate loss functions. Additionally, the comparison between (c) and NLPrompt highlights the importance of text feature initialization in our method. Among all methods, our NLPrompt achieves the best performance, with significant improvements over other baselines, further validating the effectiveness of each component.
@@ 6.7 Ablation Study
Method/Noise Ratio 30% 50% 70% Avg
@@
NLPrompt 93.44 91.21 85.38 90.01
@@ Appendix C Further Experiments
We also conducted experiments on SUN397, a dataset with a large number of classes. The results are presented in Table A6. On this dataset with a large number of classes, our NLPrompt still outperforms all baseline methods, further highlighting the superiority of our approach.
@@ Appendix E Theoretical analysis for the Robustness of PromptMAE
In this section, we demonstrate that the mean absolute error (MAE) loss is robust for prompt learning in vision-language foundation models. Leveraging the properties of vision-language pre-trained models, we assume that the latent spaces of the text encoder and image encoder are well-aligned. To clarify, we restate and explain some of our analysis settings. For a classification task, the objective is to classify an image $\mathbf{x}$ into its ground truth class $y\in[C]$, where $C$ represents the total number of classes. For simplicity, we assume that the features corresponding to these classes are orthogonal. In our theoretical analysis, we focus on a binary classification scenario, where $y_{i}\in\{+1,-1\}$.
In the definition of the text encoder (4), the incorporation of $\mathbf{W}\mathbf{p}_{c}$ introduces nonlinearity between the trainable prompt and the class prompt while maintaining the overall function’s nonlinear nature. The assumptions regarding the image encoder (6) suggest that task-relevant features differ depending on whether the label is positive or negative, while task-irrelevant features remain arbitrary and independent of the label’s polarity. Furthermore, the training loss objective is designed to strengthen the similarity between the image feature $g(\mathbf{x}_{i})$ and the text feature generated by the label class prompt $\mathbf{p}_{y_{i}}$. Following the standard feature learning theory [1], we assume that the weights of the pretrained model consist of two components: task-relevant weights $\bm{\mu}$ and task-irrelevant weights $\bm{\xi}$. We begin by proving the following feature representation lemma.
@@
Gradient analysis. This section discusses the computational approach used to analyze the performance of the algorithm, focusing on the gradient calculations essential for optimizing the model parameters. Leveraging the properties of the gradient of the Softmax function, we have:

### v2

@@ Abstract
The emergence of vision-language foundation models, such as CLIP, has revolutionized image-text representation, enabling a broad range of applications via prompt learning. Despite its promise, real-world datasets often contain noisy labels that can degrade prompt learning performance. In this paper, we demonstrate that using mean absolute error (MAE) loss in prompt learning, named PromptMAE, significantly enhances robustness against noisy labels while maintaining high accuracy. Though MAE is straightforward and recognized for its robustness, it is rarely used in noisy-label learning due to its slow convergence and poor performance outside prompt learning scenarios. To elucidate the robustness of PromptMAE, we leverage feature learning theory to show that MAE can suppress the influence of noisy samples, thereby improving the signal-to-noise ratio and enhancing overall robustness. Additionally, we introduce PromptOT, a prompt-based optimal transport data purification method to enhance the robustness further. PromptOT employs text features in vision-language models as prototypes to construct an optimal transportation matrix. This matrix effectively partitions datasets into clean and noisy subsets, allowing for the application of cross-entropy loss to the clean subset and MAE loss to the noisy subset. Our Noise-Label Prompt Learning method, named NLPrompt, offers a simple and efficient approach that leverages the expressive representations and precise alignment capabilities of vision-language models for robust prompt learning. We validate NLPrompt through extensive experiments across various noise settings, demonstrating significant performance improvements.
@@ 1 Introduction
In the realm of noisy label learning, mean absolute error (MAE) has been identified as a robust loss function within the traditional training paradigm [17]. However, MAE often suffers from slow convergence and poor performance during training, making it seldom employed as a classification loss in noise-label learning. Nevertheless, our investigation reveals an interesting phenomenon: employing MAE loss in Prompt learning (PromptMAE) notably enhances robustness while maintaining high accuracy compared to traditional cross-entropy loss. As demonstrated in Figure 1, MAE exhibits strong accuracy and fast convergence even in the presence of substantial noise.
@@
To elucidate the robustness of PromptMAE, we leverage feature learning theory [1, 3, 41], which categorizes latent representations into task-relevant and task-irrelevant components. By analyzing the optimization dynamics of these features with gradient-descent-based training, we can gain valuable insights into convergence and generalization. To this end, we find that robust prompt learning is achieved when task-relevant features dominate. Our analysis indicates that PromptMAE can suppress the influence of noisy samples, thereby enhancing robustness in prompt learning for vision-language models.
@@
In this paper, we introduce PromptOT, a prompt-based optimal transport data purification method, designed to enhance the robustness of prompt learning in vision-language foundation models. PromptOT leverages the text features as prototypes for the transportation matrix, facilitating robust prompt learning by partitioning the dataset into clean and noisy subsets. Recognizing that cross-entropy (CE) loss generally outperforms MAE on clean datasets, we apply MAE loss to train the noisy subset and cross-entropy loss to train the clean subset. This dual strategy, supported by PromptOT purification, harmonizes the strengths of both MAE and CE loss under varying noisy conditions. Our comprehensive method, named NLPrompt, leverages the expressive representation and alignment capabilities of vision-language models, offering a simple and efficient solution for robust prompt learning in the presence of noisy labels. In summary, our contributions are threefold:
@@ 2.3 Feature Learning Theory
To further understand how noisy label learning affects prompt learning, we leverage feature learning theory to analyze the learning process. Feature learning theory [1, 3, 25, 51, 23, 63, 24, 28] categorizes latent features into task-relevant and task-irrelevant components, expressing the trainable weights as a combination of these feature types. From this perspective, feature learning analyzes the coefficients of these features to gain insight into learning dynamics. Beyond its application in traditional learning paradigms, prompt learning can also be explained by feature learning theory [41]. In this paper, we adopt feature learning theory to demonstrate the robustness of MAE in prompt learning.
@@ 3 Preliminary
$\min\limits_{\mathbf{Q}\in\mathbb{R}_{+}^{n\times m}}$ $\langle\mathbf{C},\mathbf{Q}\rangle$ (1)
s.t. $\mathbf{Q}\mathds{1}_{m}=\bm{\alpha},\ \mathbf{Q}^{\top}\mathds{1}_{n}=\bm{\beta}.$
@@
s.t. $\mathbf{Q}\mathds{1}_{m}=\bm{\alpha},\ \mathbf{Q}^{\top}\mathds{1}_{n}=\bm{\beta},$
@@ 5 Methodology
Table 1: Performance metrics across various datasets and noise levels. (%)
@@ 6.3 Implementation Details
In our experiments, we adopt the same setup as CoOp [60] and JoAPR [19] to ensure a fair comparison. We use the SGD optimizer with an initial learning rate of 0.002 and employ cosine annealing. Our model backbone is consistent with CoOp [60], based on the pre-trained CLIP model [46], utilizing either ResNet-50 or ViT-B/16 as the image encoder, with ResNet-50 as the default if not explicitly specified. We use a 63M parameter text transformer as the text encoder. The default number of training epochs is set to 200. Additionally, we employ 16 shared context tokens across all categories, with the class token placed at the end. We sample a 16-shot training set from each dataset and evaluate the model on the original test set. The reported experimental results are the averages of test accuracy from three runs with different seeds, with the highest accuracy highlighted in bold. All experiments were conducted using PyTorch [43] on a cluster equipped with NVIDIA A40 GPU. It is noted that our noise setting differs from that of JoAPR. Our setting is more challenging and practical. The differences between the two noise settings and more implementation details are provided in the Supplementary Materials.
Table 2: Test accuracy (%) on Food101N.
Method CoOp GCE JoAPR NLPrompt
Accuracy 69.50 71.32 72.57 76.46
@@ 6.4 Performance Comparison
6.5 The Generalization of NLPrompt
Our method is effective not only for CoOp but also for other prompt-tuning approaches, including VPT [27], Maple [29], and PromptSRC [30], which are subsequent methods of CoOp. Additional results on the EuroSAT dataset under symmetric noise are shown in Table 3. It demonstrates the strong generalization ability of NLPrompt.
Table 3: The generalization of NLPrompt .
Method/Noise Ratio 12.5% 25.0% 37.5% 50.0% 62.5% 75.0%
VPT 89.20 79.43 65.20 61.37 41.67 27.67
VPT+Ours 91.80 91.07 89.53 86.93 80.73 73.90
MaPLe 83.27 77.67 65.27 55.40 37.53 25.47
MaPLe+Ours 89.23 84.30 78.37 76.43 73.30 59.87
PromptSRC 90.61 84.67 78.57 72.27 60.43 49.37
PromptSRC+Ours 91.29 87.67 84.97 80.33 72.10 59.50
@@ 6.6 Few-shot Learning Analysis
Method/Noise Ratio 10% 30% 50% 70% Avg
@@
NLPrompt 96.87 93.44 92.30 85.38 92.00
6.7 Ablation Study
To evaluate the effectiveness of each component of our method, we conduct ablation studies on the Flowers102 dataset. Here, we employ ViT-B/16 as the backbone and train for 100 epochs in the symmetric noise scenario. The experimental results are shown in Table 4. To validate the effectiveness of OT, we designed two sets of experiments: one without using OT for data purification and another using OT for data purification. The experimental design is as follows: (a) Use CE loss for all data; (b) Use MAE loss for all data; (c) Use random initialization prototype instead of CLIP text feature as initialization; (d) Use CE loss for clean data only after removing noisy data; (e) Use MAE loss for noisy data only after removing clean data.
The average results show that (b) outperforms (a), validating the effectiveness of our PromptMAE. Moreover, the average results show that (d) outperforms (a), and (e) outperforms (b), further validating the effectiveness of PromptOT in the data purification process. Additionally, the comparison between (c) and NLPrompt highlights the importance of text feature initialization in our method. Among all methods, our NLPrompt achieves the best performance, with significant improvements over other baselines, further validating the effectiveness of each component.
@@ 7 Conclusion
Acknowledgement
This work was supported by the National Natural Science Foundation of China (No.62406195, No.62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University) and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
@@ Appendix C Further Experiments
C.1 Experiments on SUN397
We also conducted experiments on SUN397, a dataset with a large number of classes. The results are presented in Table A6. On this dataset, our NLPrompt still outperforms all baseline methods, further highlighting the superiority of our approach.
@@ C.1 Experiments on SUN397
C.2 Experiments on Purfication Strategy
We present experiments to demonstrate that PromptOT is an effective purification method in prompt learning. We frame the purification task as a binary classification problem, where the goal is to distinguish between clean and noisy samples. To evaluate the purification performance, we use accuracy and F1-score metrics for this binary classification task. We compare our purification strategy with the partition method that uses pseudo-labels generated by CLIP zero-shot (denoted as CLIP-ZS), as well as the partition strategy used in JoAPR [19]. For our experiments, we use the Caltech101, DTD, and Flowers datasets, with results shown in Table A7. From this table, we observe that our prompt-based OT selection method achieves higher purification accuracy compared to both the CLIP zero-shot partition and the JoAPR partition.
Table A7: Comparison between different purification strategies. (%)
Method Caltech101 DTD Flowers
Acc F1 Acc F1 Acc F1
CLIP-ZS 70.38 58.35 56.25 26.72 56.62 25.16
JoAPR 50.88 44.99 51.54 32.32 50.37 32.56
NLPrompt 75.37 67.41 59.97 36.89 62.07 39.49
C.3 Hyperparameter Ablation
We evaluate the impact of different context token lengths on the DTD dataset under 50% symmetric noise in Table A8 and test various entropy regularization coefficients of OT in Figure A4. The results indicate that our NLPrompt is robust to hyperparameters.
Table A8: Test accuracy (%) under different context token lengths.
Context length 1 4 8 16 32
NLPrompt 53.65$\pm$1.95 54.00$\pm$1.49 54.07$\pm$0.40 55.20$\pm$0.33 54.73$\pm$1.52
Figure A4: Test accuracy (%) under different entropy regularization coefficients.
C.4 Comparison with JoAPR
The discrepancy in JoAPR results compared to the original paper arises from different noise settings. JoAPR introduces noisy samples uniformly across each class, whereas we randomly distribute noisy samples throughout the entire training set. For example, at a 75% noise rate with 16 shots, JoAPR’s setting turns 12 out of 16 samples in each class into noisy samples and ensuring at least 4 clean samples per class. In contrast, our setting may lead to some classes having only 1 or even no clean samples. Thus, our setting is more challenging and practical, leading to significant performance variance at high noise levels. Additionally, we implemented our NLPrompt on the DTD dataset using JoAPR’s settings and consistently observed improvements over JoAPR. The results are shown in Table A9.
Table A9: Comparison with JoAPR in the JoAPR’s noise setting.
Method/Noise Ratio 12.5% 25.0% 37.5% 50.0% 62.5% 75.0%
JoAPR 58.83 57.67 55.70 53.07 50.67 46.30
NLPrompt 63.17 61.96 60.82 59.63 53.83 49.60
C.5 Comparison with Traditional LNL Method
The traditional LNL method does not fully utilize the benefits of prompt learning with VLMs, whereas our method effectively leverage in prompt learning. In addition to the GCE method discussed in the paper, we also investigate the performance of Mixup on the EuroSAT dataset which is shown in Table A10.
Table A10: Test accuracy (%) compared with Mixup.
Method/Noise Ratio 12.5% 25.0% 37.5% 50.0% 62.5% 75.0%
CoOp 76.50 69.23 61.67 52.33 37.63 26.70
CoOp+Mixup 75.30 71.63 64.07 54.23 42.37 26.90
NLPrompt 82.53 79.53 78.13 66.70 63.53 43.80
C.6 Computational Overhead of Optimal Transport
We evaluate the computation time of optimal transport through experiments. Leveraging the efficient Sinkhorn algorithm for optimal transport, the computational overhead is minimal. For the Caltech101 dataset under a 16-shot learning setup, we perform batch matching, where 1600 image features are matched against 100 classes. The average time required for optimal transport per epoch is $0.00173$ seconds, while the original backward process takes $4.352$ seconds. Furthermore, even for significantly larger datasets, such as $100,000$ images ×1,000$\times$ 1,000× 1 , 000 classes, the average computation time for the optimal transport method is approximately $1.888$ seconds, which remains relatively small compared to the overall training time.
@@ Appendix E Theoretical Analysis for the Robustness of PromptMAE
In this section, we demonstrate that the mean absolute error (MAE) loss is robust for prompt learning in vision-language foundation models. Leveraging the properties of vision-language pre-trained models, we assume that the latent spaces of the text encoder and image encoder are well-aligned. To clarify, we restate and explain some of our analysis settings. For a classification task, the objective is to classify an image $\mathbf{x}$ into its ground truth class $y\in[C]$, where $C$ represents the total number of classes. For simplicity, we assume that the features corresponding to these classes are orthogonal. In our theoretical analysis, we focus on a binary classification scenario, where $y_{i}\in\{+1,-1\}$. In most theoretical work of feature learning, it is common to apply insights from binary classification to interpret experimental observations [1, 3]. So, we employ such theoretical frameworks to validate and support our experimental findings.
Outline In our proof, we begin by introducing the assumptions and feature modeling. We then analyze the gradient update using the chain rule and explore the relationship between the feature space and the gradient update. Next, we utilize the decomposition of trainable parameters to demonstrate how the decomposition coefficients change. By establishing the connection between the coefficients and the test loss, we compare the performance of different loss functions by examining the relationship between their respective coefficients.
Following the standard feature learning theory [1], we assume that the weights of the pretrained model consist of two components: task-relevant weights $\bm{\mu}$ and task-irrelevant weights $\bm{\xi}$. We begin by proving the following feature representation lemma.
@@
Disuccusion on feature decomposition intuition Decomposite the trainable parameters in latent space into linear combinations is a general technique in feature learning theory [1, 3]. Intuitively, in a text prompt, certain ”core” words (such as adjectives describing class features before class names) play a key role in determining the image classification and are considered ”task-relevant,” while other words are ”task-irrelevant.”
@@
Gradient analysis. In the definition of the text encoder (3), the incorporation of $\mathbf{W}\mathbf{p}_{c}$ introduces nonlinearity between the trainable prompt and the class prompt while maintaining the overall function’s nonlinear nature. The assumptions regarding the image encoder (5) suggest that task-relevant features differ depending on whether the label is positive or negative, while task-irrelevant features remain arbitrary and independent of the label’s polarity. Furthermore, the training loss objective is designed to strengthen the similarity between the image feature $g(\mathbf{x}_{i})$ and the text feature generated by the label class prompt $\mathbf{p}_{y_{i}}$. This section discusses the computational approach used to analyze the performance of the algorithm, focusing on the gradient calculations essential for optimizing the model parameters. Leveraging the properties of the gradient of the Softmax function, we have: